# Discrete State Diffusion Models: A Sample Complexity Perspective

### v1

@@ 4 Methodology
Following Zhang et al. (2025) and Chen and Ying (2024), we assume the data distribution $p_{\text{data}}$ has full support and satisfies a uniform score bound independent of the dimension $d$. Even though Zhang et al. (2025) do not explicitly assume a lower bound $1/B$, they require it for their Proposition 3 to hold. This assumption ensures well-posedness of the score function at initialization and enables dimension-free control of the score magnitude. Additionally, $\kappa_{i}=\tfrac{(p^{i}_{\text{data}})_{\max}}{(p^{i}_{\text{data}})_{\min}}$ is the ratio of the largest to smallest entry in the marginal distribution of the $i$-th dimension, measuring how unbalanced that marginal is and $\kappa^{2}=\sum_{i=1}^{d}\kappa_{i}^{2}$ aggregates this imbalance across all dimensions Lou et al. (2024) and Zhang et al. (2025) show that minimizing the discrete score entropy loss Equations (5) and (10) is equivalent to minimizing the KL divergence between the path measures of the true reverse and sampling diffusion processes, i.e., the distributions over their entire trajectories.
and by the chain rule of KL divergence,
$D_{\mathrm{KL}}(q_{0}\,\|\,p_{T})\leq D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{\pi^{d}})=D_{\mathrm{KL}}(q_{T}\,\|\,\pi^{d})+D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})$ (21)
$D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})=\frac{1}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\!\left(D_{I}\!\left(s_{t}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\right)\right)dt$ (22)
where $D_{I}\left(s_{t}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\right)$ is the Bregman divergence characterizing the distance between the true score at continuous time $s_{t}(\cdot)$ and the approximate score at discrete time points $\hat{s}_{\theta,(k+1)h}(\cdot)$, for $t\in[kh,(k+1)h]$. Here, $h$ denotes the discretization step size, and $k$ is the number of discrete time steps. Zhang et al. (2025) further decomposes Equation (22) into a score-movement term (discretization error) and a score-error term (approximation error), as given in (23), through the strong convexity of $I(\cdot)$. This decomposition facilitates analysis under their assumption of a bounded score estimation error $\epsilon_{\mathrm{score}}$. In contrast, we avoid this assumption and instead perform a rigorous analysis of the score-error term in Theorem 1 to determine the number of samples required to ensure that the overall KL divergence $D_{\mathrm{KL}}(q_{0}\,\|\,p_{T})\leq\tilde{O}(\epsilon)$, as stated in Equation (21). Using the Bregman divergence property established in Lemma 4 $(i)$, following decompostion of $D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})$ is obtained:
$D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})\;\lesssim\;$ $\tfrac{C}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\!\left[\|s_{t}(\mathbf{x}_{t})-s_{(k+1)h}(\mathbf{x}_{t})\|_{2}^{2}\right]dt$
$+\;\tfrac{C^{2}}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\!\left[D_{I}\!\left(s_{(k+1)h}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\right)\right]dt$ (23)
Earlier works such as Zhang et al. (2025) directly bound the second term on the right-hand side as $C^{2}\epsilon_{score}$. Specifically, their paper assumes access to an $\epsilon_{score}$-accurate score estimator. However, further leveraging the strong convexity of $I(\cdot)$, along with Lemma 4 $(ii)$, we further upper bound the path measure KL divergence as follows:
$D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})\;\lesssim\;\frac{C}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\bigl\|s_{t}(\mathbf{x}_{t})-s_{(k+1)h}(\mathbf{x}_{t})\bigr\|_{2}^{2}\,dt$ (24)
$\hskip 70.0001pt+\frac{C^{3}}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\bigl\|s_{(k+1)h}(\mathbf{x}_{t})-\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\bigr\|_{2}^{2}\,dt$ (25)
Note that in order to satisfy the conditions required for Lemma 4, the score functions $s_{t}(x_{t})$, $s_{(k+1)h}(x_{t})$, and $\hat{s}_{\theta,(k+1)h}(x_{t})$ must be bounded. While the true score functions are bounded through Lemma 6, the estimated score function is bounded component-wise to the range $[1/C,C]$ through hard-clipping of the final layer (Zhang et al. (2025)). From the data processing inequality (21), we obtain the following bound on $D_{\mathrm{KL}}(p_{data}\,\|\,p_{T})$ and is explained in detail in Appendix D:
$D_{KL}(p_{data}\,\|\,p_{T})\;=\ D_{KL}(q_{0}\,\|\,p_{T})\;\lesssim\;de^{-T}\log S+C\kappa^{2}S^{2}h^{2}T+\frac{M\lambda Th}{S}+\frac{C^{3}}{S}\sum_{k=0}^{K-1}hA_{k}$ (26)
The last term in (26) characterizes the score approximation error and M=C(S1)dM=C(S-1)d. To bound this error in terms of the number of samples $n_{k}$ required for training, we now consider the following loss per discrete time step. Here, for notational simplicity, we refer to the discrete time step $k$ obtained by freezing the continuous time $t$ at the upper limit of each interval. Specifically, we set, $k:=(k+1)h$, and define:
$A_{k}=\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\left\|s_{k}(\mathbf{x}_{k})\right.-\left.\hat{s}_{\theta,k}(\mathbf{x}_{k})\right\|_{2}^{2}$ (27)
In addition, we introduce a surrogate loss per discrete time step and let:
$B_{k}=\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\left\|s_{k}(\mathbf{x}_{k})\right.-\left.\overline{s}_{\theta,k}(\mathbf{x}_{k})\right\|_{2}^{2}$ (28)
Further, we introduce a clipping error:
$C_{k}=\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\left\|\bar{s}_{\theta,k}(\mathbf{x}_{k})\right.-\left.\hat{s}_{\theta,k}(\mathbf{x}_{k})\right\|_{2}^{2}$ (29)
We can decompose $A_{k}$ using the property that $(a-b)^{2}<2(a^{2}+b^{2})$ as follows:
$A_{k}\leq 2B_{k}+2C_{k}$ (30)
where we define $\overline{s}_{\theta,k}(.)$ as the unbounded score network. The term $B_{k}$ represents the difference between the true score function and the unbounded score function $\overline{s}_{\theta,k}(.)$ obtained during training. During the reverse sampling process, we use the clipped score function $\hat{s}_{\theta,k}(\mathbf{x}_{k})$. The term $C_{k}$ accounts for the error incurred due to violating the constraints of the network. We show in Lemma 13 that the clipping error $C_{k}\leq B_{k}$. This implies that
$A_{k}\leq 4B_{k}$ (31)
Now in order to upper bound $B_{k}$, we decompose it into three different components: approximation error, statistical error, and optimization error. Moreover, the term
$B_{k}\;\leq\;$ $\;\underbrace{4\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|s^{a}_{\theta,k}(\mathbf{x}_{k})-s_{k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{\mathcal{E}_{k}^{\text{approx}}}$ $+\underbrace{4\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|s^{a}_{\theta,k}(\mathbf{x}_{k})-s^{b}_{\theta,k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{\mathcal{E}_{k}^{\text{stat}}}$ $+\underbrace{4\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|\overline{s}_{\theta,k}(\mathbf{x}_{k})-s^{b}_{\theta,k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{\mathcal{E}_{k}^{\text{opt}}}$ (32)
In order to facilitate the analysis of the three terms listed above, we define the neural network parameters as follows:
$\theta_{k}^{a}=\arg\min_{\theta\in\Theta}\;\int_{kh+\delta}^{(k+1)h+\delta}\mathbb{E}_{\mathbf{x}_{t}\sim q_{t}}\left[D_{I}\left(s_{k}(\mathbf{x}_{t})\,\|\,\bar{s}_{\theta,k}(\mathbf{x}_{t})\right)\right]dt$ (33)
$\theta_{k}^{b}$ $=\arg\min_{\theta\in\Theta}\;\frac{1}{n}\sum_{i=1}^{n}D_{I}\!\left(s_{t}(\mathbf{x}_{k,t,i})\,\|\,\overline{s}_{\theta,k}(\mathbf{x}_{k,t,i})\right),t\sim\text{Unif}[kh,(k+1)h]$ (34)
where, $s^{a}_{\theta,k}$ and $s^{b}_{\theta,k}$ are the score functions associated with the parameters $\theta^{a}_{k}$ and $\theta^{b}_{k}$. Here, the loss function given in Equation (33) is the expected value of the loss function minimized at training time to obtain the unbounded estimate of the score function $\bar{s_{\theta}}$. The loss function in Equation (34) is the empirical loss function that actually is minimized to obtain $\bar{s_{\theta}}$ at training time.
We rigorously analyze each of the above errors to bound the score-approximation error in Section (6) in terms of the number of samples and neural network parameters to ensure that $D_{\mathrm{KL}}(p_{data}\,\|\,p_{T})\leq\tilde{O}(\epsilon)$.
@@ 5 Theoretical Results
Suppose Assumptions 1, 2, and 3 hold. With a probability at least $1-\delta$ the KL divergence between $p_{data}$ and $p_{T}$ is bounded by
$D_{KL}(p_{data}\,\|\,p_{T})\lesssim\;de^{-T}\log S+C\kappa^{2}S^{2}h^{2}T+\tfrac{M\lambda Th}{S}+\frac{C^{3}}{S}\sum_{k=0}^{K-1}h\,\mathcal{O}\!\left((W)^{L}\!\left((S-1)d+\tfrac{L}{W}\right)\sqrt{\tfrac{\log\!\left(\tfrac{2K}{\gamma}\right)}{n_{k}}}\;\right)$ (35)
@@
6 Proof Outline of Theorem 1
@@
In this section, we give the proof outline of the main theorem - Theorem 1 drawing upon recent advances in statistical learning theory and neural network approximation bounds. The detailed proofs are in Appendix C and D. To prove this result, we bound each of the errors in (32) and subsequently use (25) to ensure $D_{\mathrm{KL}}(p_{data}\,\|\,p_{T})\leq\tilde{O}(\epsilon)$. We use the following lemmas to bound each of these errors. The formal proofs corresponding to each of the lemmas are provided in Appendix C for completeness.
@@
Let $W$ and $D$ denote the width and depth of the neural network architecture, respectively, and we have that $W\geq(S-1)d$. Then, for all $k\in[{0},K-1]$, we have
@@
The detailed proof of this lemma is provided in Appendix C.1. We demonstrate in Lemma C.1 that if the width of the neural network $W$ is greater than the number of possible states $W\geq(S-1)d$, then it is possible to construct a neural network that is equal to any specified function defined on the (S1)d(S-1)d discrete points. The discrete nature of the score function allows us to obtain this results. In the case of diffusion models the score functions were defined on a continuous input space. In such a case, this error would have been exponential in the data dimension as is obtained in Jiao et al. (2023).
@@
We leverage assumptions 1 and 2 alongside our unique analysis to derive an upper bound on $\mathcal{E}_{\mathrm{opt}}$. The key here is our recursive analysis of the error at each stochastic gradient descent (SGD) step, which captures the error introduced by the finite number of SGD steps in estimating the score function, while prior works assumed no such error, treating the empirical loss minimizer as if it were known exactly.
@@
$\lesssim\;\underbrace{\mathcal{O}\!\left(W^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\tfrac{\log(2/\gamma)}{n_{k}}}\right)}_{\text{Statistical Error}}+\underbrace{\mathcal{O}\!\left(W^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\tfrac{\log(2/\gamma)}{n_{k}}}\right)}_{\text{Optimization Error}}$

### v2

@@ 4 Assumptions
Following Zhang et al. (2025) and Chen and Ying (2024), we assume the data distribution $p_{\text{data}}$ has full support and satisfies a uniform score bound independent of the dimension $d$. Consequently, our results in Theorem 1 do not require early stopping, and we set $\delta=0$. Even though Zhang et al. (2025) do not explicitly assume a lower bound $1/B$, they require it for their Proposition 3 to hold. This assumption ensures well-posedness of the score function at initialization and enables dimension-free control of the score magnitude. Additionally, $\kappa_{i}=\tfrac{(p^{i}_{\text{data}})_{\max}}{(p^{i}_{\text{data}})_{\min}}$ is the ratio of the largest to smallest entry in the marginal distribution of the $i$-th dimension, measuring how unbalanced that marginal is and $\kappa^{2}=\sum_{i=1}^{d}\kappa_{i}^{2}$ aggregates this imbalance across all dimensions. Lou et al. (2024) and Zhang et al. (2025) show that minimizing the discrete score entropy loss Equations (5) and (10) is equivalent to minimizing the KL divergence between the path measures of the true reverse and sampling diffusion processes, i.e., the distributions over their entire trajectories.
@@ 5 Theoretical Results
Suppose Assumptions 1, 2, and 3 hold. With a probability at least $1-\lambda$ the KL divergence between $p_{data}$ and $p_{T}$ is bounded by
$D_{KL}(p_{data}\,\|\,p_{T})\lesssim\;de^{-T}\log S+C\kappa^{2}S^{2}h^{2}T+\tfrac{M\lambda Th}{S}$
$+\frac{C^{3}}{S}\sum_{k=0}^{K-1}h\,\mathcal{O}\!\left((W)^{L}\!\left((S-1)d+\tfrac{L}{W}\right)\sqrt{\tfrac{\log\!\left(\tfrac{2K}{\gamma}\right)}{n_{k}}}\;\right)$ (20)
@@
The first term in the KL divergence bound in (20) captures the truncation error due to the insufficient mixing of the forward process. The second term reflects the discretization error due to approximating the continuous-time reverse process with a fixed number of discrete steps of size $h$. The third term is a discretization error arising from replacing the continuous-time score-error integral by its discrete-time approximation. The fourth term quantifies the approximation, statistical, and optimization errors arising from finite network capacity, finite training data, and imperfect minimization of the empirical loss.
@@
6 Proof of Theorem 1
and by the chain rule of KL divergence,
$D_{\mathrm{KL}}(q_{0}\,\|\,p_{T})\leq D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{\pi^{d}})$
$=D_{\mathrm{KL}}(q_{T}\,\|\,\pi^{d})+D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})$ (23)
$D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})$
$=\frac{1}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{x_{t}}\!\left(D_{I}\!\left(s_{t}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\right)\right)dt$ (24)
where the expectation is taken over $\mathbf{x}_{t}\sim q_{t}$} and $D_{I}\left(s_{t}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\right)$ is the Bregman divergence characterizing the distance between the true score at continuous time $s_{t}(\cdot)$ and the approximate score at discrete time points $\hat{s}_{\theta,(k+1)h}(\cdot)$, for $t\in[kh,(k+1)h]$. Here, $h$ denotes the discretization step size, and $k$ is the number of discrete time steps. Zhang et al. (2025) further decomposes Equation (24) into a score-movement term (discretization error) and a score-error term (approximation error), as given in (25), through the strong convexity of $I(\cdot)$. This decomposition facilitates analysis under their assumption of a bounded score estimation error $\epsilon_{\mathrm{score}}$. In contrast, we avoid this assumption and instead perform a rigorous analysis of the score-error term in Theorem 1 to determine the number of samples required to ensure that the overall KL divergence $D_{\mathrm{KL}}(q_{0}\,\|\,p_{T})\leq\tilde{O}(\epsilon)$, as stated in Equation (23). Using the Bregman divergence property established in Lemma 4 $(i)$, following decompostion of $D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})$ is obtained:
$D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})$
$\lesssim\tfrac{C^{2}}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}\,D_{I}\!\left(s_{(k+1)h}(\mathbf{x}_{t})\,\|\,\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\right)dt$
$+\tfrac{C}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}\,\|s_{t}(\mathbf{x}_{t})-s_{(k+1)h}(\mathbf{x}_{t})\|_{2}^{2}\,dt$ (25)
where the expectations are taken over $\mathbf{x}_{t}\sim q_{t}$}. Earlier works such as Zhang et al. (2025) directly bound the first term on the right-hand side as $C^{2}\epsilon_{score}$. Specifically, their paper assumes access to an $\epsilon_{score}$-accurate score estimator. The notation $y\lesssim z$ means that there exists a universal constant $C^{\prime}>0$ such that $y\leq C^{\prime}z$. However, further leveraging the strong convexity of $I(\cdot)$, along with Lemma 4 $(ii)$, we further upper bound the path measure KL divergence in (25) as follows:
$D_{\mathrm{KL}}(\mathbb{Q}\,\|\,\mathbb{P}^{q_{T}})\lesssim\;$
$\ \frac{C}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{x_{t}}\bigl\|s_{t}(\mathbf{x}_{t})-s_{(k+1)h}(\mathbf{x}_{t})\bigr\|_{2}^{2}\,dt+$
$\frac{C^{3}}{S}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{x}_{t}}\bigl\|s_{(k+1)h}(\mathbf{x}_{t})-\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{t})\bigr\|_{2}^{2}\,dt$ (26)
where the expectations are taken over $\mathbf{x}_{t}\sim q_{t}$. Note that in order to satisfy the conditions required for Lemma 4, the score functions $s_{t}(x_{t})$, $s_{(k+1)h}(x_{t})$, and $\hat{s}_{\theta,(k+1)h}(x_{t})$ must be bounded. While the true score functions are bounded through Lemma 6, the estimated score function is bounded component-wise to the range $[1/C,C]$ through hard-clipping of the final layer (Zhang et al., 2025). From the data processing inequality (23), we obtain the following bound on $D_{\mathrm{KL}}(p_{data}\,\|\,p_{T})$ and is explained in detail in Appendix D:
$D_{KL}(p_{data}\,\|\,p_{T})\;=\ D_{KL}(q_{0}\,\|\,p_{T})$
$\;\lesssim\;de^{-T}\log S+C\kappa^{2}S^{2}h^{2}T+\frac{M\lambda Th}{S}+\frac{C^{3}}{S}\sum_{k=0}^{K-1}hA_{k}$ (27)
The last term in (27) characterizes the score approximation error and M=C(S1)dM=C(S-1)d. To bound this error in terms of the number of samples $n_{k}$ required for training, we now consider the following loss per discrete time step. Here, for notational simplicity, we refer to the discrete time step $k$ obtained by freezing the continuous time $t$ at the upper limit of each interval. Specifically, we write, $k:=(k+1)h$. To facilitate our analysis, we introduce the unbounded ($\overline{s}_{\theta,k}(\cdot)$) and the clipped ($\hat{s}_{\theta,k}(\mathbf{x}_{k})$) score functions. During the reverse sampling process, we use the clipped score function $\hat{s}_{\theta,k}(\mathbf{x}_{k})$ by projecting $\overline{s}_{\theta,k}(\cdot)$ into the interval $[1/C,C]$ component-wise. Using these definitions, we obtain the following loss per discrete time step $k$:
$\leq\underbrace{2\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|s_{k}(\mathbf{x}_{k})-\bar{s}_{\theta,k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{B_{k}}\quad+$
$\quad\underbrace{2\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|\bar{s}_{\theta,k}(\mathbf{x}_{k})-\hat{s}_{\theta,k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{C_{k}}$ (28)
where the inequality in (28) follows from $(a+b)^{2}\leq 2(a^{2}+b^{2})$ by writing $A_{k}=\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\Bigl\|\bigl(s_{k}(\mathbf{x}_{k})-\bar{s}_{\theta,k}(\mathbf{x}_{k})\bigr)+\bigl(\bar{s}_{\theta,k}(\mathbf{x}_{k})-\hat{s}_{\theta,k}(\mathbf{x}_{k})\bigr)\Bigr\|_{2}^{2}$.
The term $B_{k}$ represents the difference between the true score function and the unbounded score function $\overline{s}_{\theta,k}(\cdot)$ obtained during training. The term $C_{k}$ accounts for the error incurred due to violating the constraints of the network. We show in Lemma 13 that the clipping error $C_{k}\leq B_{k}$. This implies that
$A_{k}\leq 4B_{k}$ (29)
Now in order to upper bound $B_{k}$, we decompose it into three different components: approximation error, statistical error, and optimization error. In order to facilitate the analysis of the three terms listed above, we define the following neural network parameters.
$\theta_{k}^{a}=\arg\min_{\theta\in\Theta}\;\int_{kh}^{(k+1)h}\mathbb{E}_{x_{t}}\left[D_{I}\left(s_{k}(\mathbf{x}_{t})\,\|\,\bar{s}_{\theta,k}(\mathbf{x}_{t})\right)\right]dt$ (30)
$\theta_{k}^{b}$ $=\arg\min_{\theta\in\Theta}\;\frac{1}{n}\sum_{i=1}^{n}D_{I}\!\left(s_{t}(\mathbf{x}_{k,t,i})\,\|\,\overline{s}_{\theta,k}(\mathbf{x}_{k,t,i})\right)$ (31)
$\quad t\sim\text{Unif}[kh,(k+1)h]$
Moreover, the term
$B_{k}\;\leq\;$ $\;\underbrace{4\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|s^{a}_{\theta,k}(\mathbf{x}_{k})-s_{k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{\mathcal{E}_{k}^{\text{approx}}}$
$+\underbrace{4\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|s^{a}_{\theta,k}(\mathbf{x}_{k})-s^{b}_{\theta,k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{\mathcal{E}_{k}^{\text{stat}}}$
$+\underbrace{4\,\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\bigl\|\overline{s}_{\theta,k}(\mathbf{x}_{k})-s^{b}_{\theta,k}(\mathbf{x}_{k})\bigr\|_{2}^{2}}_{\mathcal{E}_{k}^{\text{opt}}}$ (32)
where, $s^{a}_{\theta,k}(\cdot)$ and $s^{b}_{\theta,k}(\cdot)$ are the score functions associated with the parameters $\theta^{a}_{k}$ and $\theta^{b}_{k}$. Here, the loss function given in Equation (30) is the expected value of the loss function minimized at training time to obtain the unbounded estimate of the score function $\overline{s}_{\theta}(\cdot)$. The loss function in Equation (31) is the empirical loss function that actually is minimized to obtain $\overline{s}_{\theta}(\cdot)$ at training time.
@@ 6 Proof of Theorem 1
Drawing upon recent advances in statistical learning theory and neural network approximation bounds, we bound each of the above errors through Lemmas 1, 2, and 3 to bound the score-estimation error in terms of the number of samples and neural network parameters. The detailed proofs are provided in Appendix C and D. To prove Theorem 1, we bound each of the errors in (32) and subsequently use (26) to ensure $D_{\mathrm{KL}}(p_{data}\,\|\,p_{T})\leq\tilde{O}(\epsilon)$. We introduce the following lemmas to bound the approximation, statistical, and optimization errors. The formal proofs corresponding to each of the lemmas are provided in Appendix C for completeness.
@@
Let $W$ and $L$ denote the width and depth of the neural network architecture, respectively, and we have that $W\geq(S-1)d$. Then, for all $k\in[{0},K-1]$, we have
@@
The detailed proof of this lemma is provided in Appendix C.1. We demonstrate in Lemma C.1 that if the width of the neural network $W$ is greater than the number of possible states $W\geq(S-1)d$, then it is possible to construct a neural network that is equal to any specified function defined on the (S1)d(S-1)d discrete points, where width $W$ is defined as the maximum number of neurons in a layer of the neural network. The discrete nature of the score function allows us to obtain this results. In the case of diffusion models the score functions were defined on a continuous input space. In such a case, this error would have been exponential in the data dimension as is obtained in Jiao et al. (2023).
@@
We leverage assumptions 1 and 2 alongside our unique analysis to derive an upper bound on $\mathcal{E}_{\mathrm{opt}}$. The key here is our recursive analysis of the error at each stochastic gradient descent (SGD) step, which captures the error introduced by the finite number of SGD steps in estimating the score function, while prior works assumed no such error, treating the empirical loss minimizer as if it were known exactly. Here, the dependence on the number of SGD steps is implicit through $n_{k}$, since each SGD update at step $k$ uses a fresh sample and the total number of optimization iterations equals the effective sample size $n_{k}$.
@@
$\lesssim\;\underbrace{\mathcal{O}\!\left(W^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\tfrac{\log(2/\gamma)}{n_{k}}}\right)}_{\text{Statistical Error}}$ (36)
$+\underbrace{\mathcal{O}\!\left(W^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\tfrac{\log(2/\gamma)}{n_{k}}}\right)}_{\text{Optimization Error}}$